OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form, and find the center and radius of the circle. x2 + y2 − 8x + 12y + 27 = 0

OpenStudy (anonymous):

\[(x-4)^2+(y+6)^2=5^2\] center(4,-6) ; Radius=5